OpenStudy (quickstudent):

Can someone simplify square root of 44? I don't remember how to do it, so I need an explanation of the steps.

OpenStudy (anonymous):

what are your answer choices?

OpenStudy (quickstudent):

There aren't any choices, I have to write it myself.

OpenStudy (anonymous):

i got 6.63 on my calculator

OpenStudy (quickstudent):

But I need to have the root simplified, an answer like, maybe\[11 \sqrt{5}\]

OpenStudy (quickstudent):

That's my example, not guess

OpenStudy (quickstudent):

Wait a minute, I think I just remembered how to do it. \[2^{2}*11\] \[\sqrt{2^2}*\sqrt{11}\] \[2\sqrt{11}\] Is this correct?